# Aging-aware Energy Management for Residential Multi-Carrier Energy Systems

### v2

@@ 7 Acknowledgment
The project was carried out with a Top Sector Energy subsidy from the Ministry of Economic Affairs and Climate, carried out by the Netherlands Enterprise Agency (RVO). The specific subsidy for this project concerns the MOOI subsidy round 2020.
@@ Appendix A Appendix: A
$OCV_{n,b,t}(z_{b,t})$ $=0.6379+0.5416\cdot e^{-305.5309\cdot z_{b,t}}$ (47)
$\quad+0.044\cdot\tanh\left(-\frac{z_{b,t}-0.1958}{0.108}\right)$
$\quad-0.1978\cdot\tanh\left(\frac{z_{b,t}-1.0571}{0.0854}\right)$
$\quad-0.6875\cdot\tanh\left(\frac{z_{b,t}+0.0117}{0.0529}\right)$
$\quad-0.0175\cdot\tanh\left(\frac{z_{b,t}-0.5692}{0.0875}\right)$
@@
Parameter Description Units NMC (SANYO cell) LFP (A123 cell) $n_{\textrm{SEI}}$ Number of e- transferred in SEI side reaction - 2.0 $\lambda$ Constant $\lambda=\frac{c_{s}\sqrt{D_{s}}}{c_{p}\sqrt{D_{p}}}$ 5.51×$\times 10^{-5}$ $OCV_{s}$ OCV of the side reaction V 0.4 $\varepsilon_{\textrm{AM}}$ Active material volume fraction - 0.552 $R_{s}$ Particle radius m 7.5×$\times 10^{-6}$ 5×$\times 10^{-6}$ $a_{s}$ Specific surface area of the anode m-1 $\frac{3\cdot\varepsilon_{\textrm{AM}}}{R_{s}}$ $A_{n}$ Active surface area of the anode m2 0.105 0.18 $L_{n}$ Thickness of anode m 50×$\times 10^{-6}$ 34×$\times 10^{-6}$ $i_{0}$ Exchange current of the intercalation current A/m2 1.5 $k_{\textrm{SEI}}$ Kinetic rate 1/$\sqrt{\text{s}}$ 66.85 $E_{\textrm{SEI}}$ Activation energy J/mol 39146.0 $\delta_{SEI,0}$ Initial value of the SEI layer thickness m 2.0 ×$\times 10^{-9}$ $M_{\textrm{SEI}}$ Molecular weight of the SEI layer kg/mol 0.162 $\rho_{\textrm{SEI}}$ Density of the SEI layer kg/m3 1690.0 $z_{100\%}$ Full electrode stoichiometry - 0.9 0.81 $z_{0\%}$ Empty electrode stoichiometry - 0 0.0176 $k_{\textrm{AM}}$ Kinetic rate 1/Ah 0.0137 $E_{\textrm{AM}}$ Activation energy J/mol 39500.0 $\beta$ Tuning parameter - 1.7 $t^{+}_{0}$ Transport/transference number - 0.363 0.36 $E_{\kappa}$ Activation energy for $\kappa$ J/mol 34700.0 $E_{D_{e}}$ Activation energy for $D_{e}$ J/mol 34700.0 $\kappa_{\textrm{ref}}$ Reference ionic conductivity for $\kappa$ at reference temperature S/m 0.174 $D_{e,\textrm{ref}}$ Reference value for $D_{e}$ at reference temperature m2/s 7.5×$\times 10^{-11}$ $brug$ Bruggeman coefficient - 3/2 $c_{e,\textrm{avg}}$ Average volume concentration of Li in the electrolyte mol/m3 1000 1200 $c_{e,\textrm{max}}$ Maximum volume concentration of Li in the electrolyte mol/m3 1000 1200 $\sigma_{n}$ Electronic conductivity S/m 100 215 $\varepsilon_{s}$ Volume fraction of solid electrolyte - 0.59 0.58
@@
Glossary
Acronyms
AM
active material loss
BESS
battery energy storage system
BM
bucket model
BMS
battery management system
DER
distributed energy resources
DLA
Direct Lookahead
ECM
equivalent circuit model
EMA
energy management algorithm
EMS
energy management systems
ESS
energy storage systems
EV
electric vehicle
FOM
full order model
HESS
hybrid energy storage system
HP
heat pump
LFP
Lithium iron phosphate
MCES
multicarrier energy systems
MPC
Model Predictive Control
NMC
nickel manganese cobalt oxides
OCP
optimal control problem
PB
physics-based
PBROM
physics-based reduced order model
PEI
power electronic interface
RL
reinforcement learning
SCM
series connected module
SPD
sequential decision problem
SEI
solid electrolyte interface
SPV
solar photovoltaics
ST
solar thermal
TESS
thermal energy storage system
UMF
Universal Modeling Framework

### v3

@@ 7 Acknowledgment
The project was carried out with a Top Sector Energy subsidy from the Ministry of Economic Affairs and Climate, carried out by the Netherlands Enterprise Agency (RVO). The specific subsidy for this project concerns the MOOI subsidy round 2020, FLEXINet project grant number MOOI 32027.
8 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author(s) used ChatGPT and Gemini to help the writing process in favor of clarity and readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@ Appendix A Appendix: A
[…]
@@
[…]